A) calculate the ph of a 0.028 M HNO3 solution
B) Calculate the pH of a 0.028 M solution of benzoic acid (Ka=6.5x10^-5)
Answer #1

A- The given acid HNO3 is a strong acid. We know the strong acids dissociates 100% to H+ ion. Thus the pH of strong acid is calculated as

pH = -log [H+]

Now due to 100% dissociation i.e all the HNO3 will completely converted into H+, [H+] = [HNO3] =  0.028 M

Thus pH = -log [H+] =  -log [0.028] = 1.55

B-

For Benzoic acid (C6H5COOH), it is a weak acid. Thus it cant dissociates 100% to H+ ion. Hence some of the acid dissociate ti H+ and some remain as C6H5COOH.

Now the given Ka represents the dissociation constant for C6H5COOH

i.e C6H5COOH → H+ + C6H5COO- This is the dissociation reaction of Benzoic acid

Here  C6H5COOH = the weak acid and C6H5COO- = conjugate base

Here Ka = [H+][C6H5COO-]/[C6H5COOH]

Now pH of weak acid is calculated by Henderson Hasselblach Equation-

pH = pKa + log [Conjugate base] / [weak acid] where pKa = -log Ka

Now given [C6H5COOH] = 0.028 M and Ka = 6.5x10-5

Lets consider x M of acid is dissociate to C6H5COO- . Then to find the concentration of acicd and conjugate base, we have to form the ICE table

[C6H5COOH] [H+] [C6H5COO-]
Initial 0.028 M 0 0
Change -x +x +x
Equilibrium 0.028 M -x +x +x

Now putting the values for

Ka = [H+][C6H5COO-]/[C6H5COOH]

6.5x10-5 = x * x / 0.028 M -x

6.5x10-5 * (0.028 M -x) = x2

0.00000182‬ - 0.000065x = x2

Or x2‬ + 0.000065x - 0.00000182 = 0

Now solving this,

x = 0.0013

Thus at equilibrium, [C6H5COOH] = 0.028 M -x = 0.028 M - 0.0013 = 0.0267‬ M

And [C6H5COO-] = x =  0.0013 M

Again pKa = -log Ka = -log (6.5x10-5) = 5 - log 6.5 = 4.18

Now putting the values in the equation-

pH = pKa + log [Conjugate base] / [weak acid]

pH = 4.18 + log [0.0013] / [0.0267‬]

= 4.18 + log [0.048‬]

= 2.86